How to determine the H3O+ of the buffer solution made by mixing two solutions? Given volume and concentration for both and the pKa.

For example, mixing 6.4 mL of 1.3 M acetic acid and 3.6 mL of 0.85 M sodium acetate. The pKa for the acetic acid being 4.8.
Answer #1

The number of millimoles of acetic acid = molarity* volume in mL = 1.3*6.4 = 8.32 mmoles
The number of millimoles of sodium acetate        = 0.85*3.6 = 3.06 mmoles

pH of acidic buffer = pKa + log(CH3COONa/ CH3COOH)

                              = 4.8 + log(3.06/8.32) = 4.37

But, pH = -log(H3O+) = 4.37

Therefore, the concentration of H3O+ ion, [H3O+] = 4.27*10^-5 M